OpenStudy (anonymous):

3/48 is what part of 1/12

OpenStudy (anonymous):

3/4 th part (1/12)*k = 3/48 k=(3/48)*12=3/4

OpenStudy (anonymous):

thankx

OpenStudy (anonymous):

A simple way to remember this is that 4/48 is 1/12. 3/48 is one less 3/4